OpenStudy (anonymous):

integrate x^2/ sqrt(4+x^2)

hartnn (hartnn):

did you try u= 4+x^2 ?

OpenStudy (anonymous):

no, let me do that

hartnn (hartnn):

cool, let me know if you get stuck on any step :)

hartnn (hartnn):

another way i could think of is to write x^2 as x^2+4-4 and then splitting the fraction sqrt (x^2+4)-4/sqrt(x^2+4)

OpenStudy (anonymous):

du=2xdx and im not sure what to do with the x

hartnn (hartnn):

x= \sqrt (u-4) now find dx

OpenStudy (anonymous):

my new integral looks like \[\int\limits (u-4)/ 2(\sqrt{u \sqrt{u-4}})\]

OpenStudy (xapproachesinfinity):

xdx=du you don't need to do anything you x^2 on top make it x.xdx

OpenStudy (xapproachesinfinity):

However, this is little different from what @hartnn suggested

OpenStudy (anonymous):

so what would u be?

OpenStudy (anonymous):

Some other suggestions for substituting: try \(x=2\tan u\) or \(x=\sqrt t\).

OpenStudy (xapproachesinfinity):

oh yeah i like that sub @SithsAndGiggles suggested just now makes thing a lot simpler try it^_^

OpenStudy (xapproachesinfinity):

i meant the x=2tanu

OpenStudy (xapproachesinfinity):

this requires some trig identities to do^_^

OpenStudy (anonymous):

how do i make x=2tanu?

OpenStudy (xapproachesinfinity):

for u as you started first u=4+x^2 du=2xdx =====> 1/2du=xdx if you want to continue this way

OpenStudy (xapproachesinfinity):

set x=2tanu then x^2=4tan^2u

OpenStudy (xapproachesinfinity):